Need to answer in a+bi form where a and b are real numbers. (5-6i)+(3-4i)
Found 2 solutions by ikleyn, Edwin McCravy:
Answer by ikleyn(52778)


(5-6i) + (3-4i) = (5+3) + (-6i-4i) = 8 + (-10)i.


Answer by Edwin McCravy(20054)

(5-6i)+(3-4i)

Remove the parentheses

5-6i+3-4i

Add the like terms 5 and 3 and get 8
Add the like terms -6i and -4i and get -10i

8-10i   Easy, huh?

Edwin